# Topological constraints on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v2

@@ 2.5.2 Exact Lagrangian cobordism arising from clean intersection
Let $L$ and $L^{\prime}$ be submanifolds of $T^{*}\mathbb{R}^{3}$. We say that $L$ and $L^{\prime}$ have a clean intersection along $L\cap L^{\prime}$ if the following hold:
  • $L\cap L^{\prime}$ is a submanifold of both $L$ and $L^{\prime}$.
  • For every $x\in L\cap L^{\prime}$, $T_{x}(L\cap L^{\prime})=T_{x}L\cap T_{x}L^{\prime}$.
Let us make a setup to deal with Question 1.1. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and let $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$. Suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. By moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$, we may assume that there exists $\varepsilon>0$ such that
$\varphi(L_{K_{0}})\cap D_{\varepsilon}^{*}\mathbb{R}^{3}=L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$
For the proof, see [16, Lemma 2.3]. We consider a diffeomorphism
@@
for which $F^{*}(\lambda_{\mathbb{R}^{3}})=e^{r}\alpha_{\mathbb{R}^{3}}$. Furthermore,
$L_{\varphi}\coloneqq F^{-1}(\varphi(L_{K_{0}})\setminus K_{1})$
is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. We have the inclusion maps (1)
@@
The proof is left to Appendix A. Note that the first assertion about orientations was shown in [16, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [16, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ together with the triangle product. It should be possible to prove Proposition 2.7 in a way parallel to [16, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$. However, we will give a simpler proof in Appendix A by using Floer cohomology with coefficients in $\mathbb{F}_{2}[H_{1}(L_{K_{0}})]\cong\mathbb{F}_{2}[\lambda^{\pm}]$.
@@
It is clear that $(i_{\pm})_{*}\colon R_{\pm}\to R_{0}$ maps $U$ to $U_{0}$. The circle $\varphi(K_{0})$ in $\varphi(L_{K_{0}})$ is oriented via the diffeomorphism $\left.\varphi\right|_{K_{0}}\colon K_{0}\to\varphi(K_{0})$. From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $K_{1}]=a\cdot[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$. Consider the map induced by $i_{\pm}$ on the first homology groups
$\begin{array}[]{ll}(i_{+})_{*}\colon H_{1}(\Lambda_{K_{0}})\to H_{1}(L_{\varphi}),&(i_{-})_{*}\colon H_{1}(\Lambda_{K_{1}})\to H_{1}(L_{\varphi}).\end{array}$
A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that the equations
@@
Next, take an embedded disk $D\subset\varphi(L_{K_{0}})$ which bounds the circle $(F\circ i_{+}\circ m)(S^{1})$. $D$ is oriented so that
[…]
with respect to the diffeomorphism $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$. We may assume that $D$ intersects $K_{1}$ transversely. Then, their algebraic intersection number is equal to $a\in\{\pm 1\}$. For sufficiently small $\varepsilon>0$, $D\cap(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})=\coprod_{i\in I}D^{\prime}_{i}$, where $D^{\prime}_{i}\subset D$ is a small disk and $I$ is a finite set. The summand of $\partial D_{i}]\in H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$ for all $i\in I$ satisfy
@@
$L_{K_{0}}$ is equipped with the spin structure $\mathfrak{s}_{L_{K_{0}}}$ given by a trivial $\mathrm{Spin}(3)$-bundle. We restrict $\mathfrak{s}_{L_{K_{0}}}$ on $L_{K_{0}}\setminus\varphi^{-1}(K_{1})$. Then via the diffeomorphism $F^{-1}\circ\varphi\colon L_{K_{0}}\setminus\varphi^{-1}(K_{1})\to L_{\varphi}$, we obtain a spin structure $s_{L_{\varphi}}$ on the exact Lagrangian cobordism $L_{\varphi}$. This induces $\mathfrak{s}_{K_{0}}$ on the positive end $\Lambda_{K_{0}}$ and $\mathfrak{s}_{K_{1}}$ on the negative end $\Lambda_{K_{0}}$. (Here, we need to care about the underlying orientations of spin manifolds. See the arguments in [16, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6. ∎
@@ 3.1 General case
We will combine Proposition 2.8 with the results in [3, 14, 15] about the knot DGA. Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [15, Appendix A]. In this section, we mainly refer to [15] for the definition of the knot DGA.
Let $K$ be an oriented knot in $\mathbb{R}^{3}$ given by the closure of a braid $B$. Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [15, Definition 3.11]. Its stable tame isomorphism class as a DGA over $R$ depends only on the isotopy class of $K$ as an oriented knot. See [14, Theorem 1.4] and [15, Theorem 3.1]. Let us denote this DGA by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ and its homology $\operatorname{Ker}\partial_{K}/\operatorname{Im}\partial_{K}$ by $HC_{*}(K)$. We will review a description of the 0-th degree part $HC_{0}(K)$ in Section 3.2.
@@
The base change by $\psi^{\prime}$ is due to the difference of conventions between [3] and [15]. See [15, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
@@
We continue to consider the setup in Subsection 2.5.2. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$. For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. Then we have the following theorem analogous to [16, Proposition 4.6].
@@
We recall the definition of augmentations of a DGA. See, for instance, [15, Definition 2.6].
@@
Let us prepare a lemma which is necessary to deduce Proposition 3.10 and Proposition 3.17.
Lemma 3.4.
Let $(\mathcal{A}_{*},\partial)$ be a unital non-commutative DGA over $R$ and let $\mathbb{k}$ be a field. Let $H^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial)$ denote the commutative $R$-algebra obtained by the abelianization of the 0-th homology of $(\mathcal{A}_{*},\partial)$. Suppose that there exist a polynomial $P\in R[T]$ and a unital $R$-algebra map
$J\colon R[T]/(P)\to H^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial).$
Then, for any augmentation $\varepsilon$ of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)\in(\mathbb{k}^{*})^{3}$, the polynomial $\left.P\right|_{\lambda=x,\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
@@
Any augmentation $\varepsilon$ of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$ induces a unital ring homomorphism $\bar{\varepsilon}\colon HC^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial)\to\mathbb{k}$ such that $(\bar{\varepsilon}(\lambda),\bar{\varepsilon}(\mu),\bar{\varepsilon}(U))=(x,y,Z)$. We obtain a unital ring homomorphism $\bar{\varepsilon}\circ J\colon R[T]/(P)\to\mathbb{k}$, and $(\bar{\varepsilon}\circ J)(T)\in\mathbb{k}$ is a root of $\left.P\right|_{\lambda=x,\mu=y,U=Z}\in\mathbb{k}[T]$. ∎
The next proposition is used to prove Theorem 3.15 and Theorem 3.19 by taking $\mathbb{k}=\mathbb{Q}$.
@@
Suppose that $K_{1}$ is the unknot. Then, for any $y,Z\in\mathbb{k}^{*}$ satisfying $y\neq Z^{n}$ for every $n\in\mathbb{Z}$, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)$.
@@
If $K_{1}$ is the unknot, from the trivial braid on a single strand, we can compute as in [15, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by
$U-\lambda-\mu+\lambda\mu.$
Let us take $a\in\{\pm 1\}$ and $b,c,d\in\mathbb{Z}$ to be the integers in Theorem 3.2. For any $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, we take an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{k}$ determined by
@@
Note that $y\cdot Z^{-d}\neq 1$ and $y\cdot Z^{-d}\neq Z^{a}$ from the condition on $y$ and $Z$.
Composing with the map $\Psi_{\varphi}$ of Theorem 3.2, we obtain $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathbb{k}$. It is an augmentation of the knot DGA of $K_{0}$ to $\mathbb{k}$ such that
$\begin{array}[]{cc}\varepsilon(\mu)=\varepsilon_{1}(\mu^{a}U^{d})=y,&\varepsilon(U)=\varepsilon_{1}(U)=Z.\end{array}$
Therefore, $\varepsilon$ satisfies the required condition. ∎
Let us also consider a knot invariant defined from $\mathbb{C}$-valued augmentations of the knot DGA.
Definition 3.6 (Definition 5.1 of [15]).
For an oriented knot $K$ in $\mathbb{R}^{3}$, its augmentation variety is a subset of $(\mathbb{C}^{*})^{3}$ defined by
[…]
From Theorem 3.2, we obtain a refinement of [16, Theorem 1.1].
Corollary 3.7.
There exist $a\in\{\pm 1\}$ and integers $b,c$ and $d^{\prime}$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},yZ^{d},Z).$
Proof.
Using $\Psi_{\varphi}$ of Theorem 3.2, it follows that for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{C}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{C}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{C}$ such that
From the definition of $V_{K_{i}}$ for $i\in\{0,1\}$, we obtain a map from $V_{K_{1}}$ to $V_{K_{0}}$ as in the assertion. ∎

### v3

@@ 2.5.2 Exact Lagrangian cobordism arising from clean intersection
Let $L$ and $L^{\prime}$ be submanifolds of $T^{*}\mathbb{R}^{3}$. We say that $L$ and $L^{\prime}$ have a clean intersection along $L\cap L^{\prime}$ if $L\cap L^{\prime}$ is a submanifold of both $L$ and $L^{\prime}$, and $T_{x}(L\cap L^{\prime})=T_{x}L\cap T_{x}L^{\prime}$ for every $x\in L\cap L^{\prime}$.
Let us make a setup. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and let $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$. Suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. By moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$, we may assume that there exists $\varepsilon>0$ such that $\varphi(L_{K_{0}})\cap D_{\varepsilon}^{*}\mathbb{R}^{3}=L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$. For the proof, see [15, Lemma 2.3]. We consider a diffeomorphism
@@
for which $F^{*}(\lambda_{\mathbb{R}^{3}})=e^{r}\alpha_{\mathbb{R}^{3}}$. Furthermore, $L_{\varphi}\coloneqq F^{-1}(\varphi(L_{K_{0}})\setminus K_{1})$ is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. We have the inclusion maps (2)
@@
The proof is left to Appendix A. Note that the first assertion about orientations was shown in [15, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [15, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ together with the triangle product. It should be possible to prove Proposition 2.7 in a way parallel to [15, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$. However, we will give a simpler proof in Appendix A by using Floer cohomology with coefficients in $\mathbb{F}_{2}[H_{1}(L_{K_{0}})]\cong\mathbb{F}_{2}[\lambda^{\pm}]$.
@@
It is clear that $(i_{\pm})_{*}\colon R_{\pm}\to R_{0}$ maps $U$ to $U_{0}$. The circle $\varphi(K_{0})$ in $\varphi(L_{K_{0}})$ is oriented via the diffeomorphism $\left.\varphi\right|_{K_{0}}\colon K_{0}\to\varphi(K_{0})$. From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $K_{1}]=a\cdot[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$. Consider the map induced by $i_{\pm}$ on the first homology groups $(i_{+})_{*}\colon H_{1}(\Lambda_{K_{0}})\to H_{1}(L_{\varphi})$ and $(i_{-})_{*}\colon H_{1}(\Lambda_{K_{1}})\to H_{1}(L_{\varphi})$. A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that the equations
@@
Next, take an embedded disk $D\subset\varphi(L_{K_{0}})$ which bounds the circle $(F\circ i_{+}\circ m)(S^{1})$. $D$ is oriented so that $\partial D]=(\left.F\right|_{L_{\varphi}})_{*}((i_{+})_{*}([m]))$ in $H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$ with respect to the diffeomorphism $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$. We may assume that $D$ intersects $K_{1}$ transversely. Then, their algebraic intersection number is equal to $a\in\{\pm 1\}$. For sufficiently small $\varepsilon>0$, $D\cap(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})=\coprod_{i\in I}D^{\prime}_{i}$, where $D^{\prime}_{i}\subset D$ is a small disk and $I$ is a finite set. The summand of $\partial D_{i}]\in H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$ for all $i\in I$ satisfy
@@
$L_{K_{0}}$ is equipped with the spin structure $\mathfrak{s}_{L_{K_{0}}}$ given by a trivial $\mathrm{Spin}(3)$-bundle. We restrict $\mathfrak{s}_{L_{K_{0}}}$ on $L_{K_{0}}\setminus\varphi^{-1}(K_{1})$. Then via the diffeomorphism $F^{-1}\circ\varphi\colon L_{K_{0}}\setminus\varphi^{-1}(K_{1})\to L_{\varphi}$, we obtain a spin structure $s_{L_{\varphi}}$ on the exact Lagrangian cobordism $L_{\varphi}$. This induces $\mathfrak{s}_{K_{0}}$ on the positive end $\Lambda_{K_{0}}$ and $\mathfrak{s}_{K_{1}}$ on the negative end $\Lambda_{K_{0}}$. (Here, we need to care about the underlying orientations of spin manifolds. See the arguments in [15, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6. ∎
@@ 3.1 General case
We will combine Proposition 2.8 with the results in [3, 13, 14] about the knot DGA. Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [14, Appendix A]. In this section, we mainly refer to [14] for the definition of the knot DGA.
Let $K$ be an oriented knot in $\mathbb{R}^{3}$ given by the closure of a braid $B$. Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [14, Definition 3.11]. Its stable tame isomorphism class as a DGA over $R$ depends only on the isotopy class of $K$ as an oriented knot. See [13, Theorem 1.4] and [14, Theorem 3.1]. Let us denote this DGA by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ and its homology $\operatorname{Ker}\partial_{K}/\operatorname{Im}\partial_{K}$ by $HC_{*}(K)$. We will review a description of the 0-th degree part $HC_{0}(K)$ in Section 3.2.
@@
The base change by $\psi^{\prime}$ is due to the difference of conventions between [3] and [14]. See [14, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
@@
We continue to consider the setup in Subsection 2.5.2. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$. For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. Then we have the following theorem analogous to [15, Proposition 4.6].
@@
We recall the definition of augmentations of a DGA. See, for instance, [14, Definition 2.6].
@@
Let us also consider a knot invariant defined from $\mathbb{k}$-valued augmentations of the knot DGA.
Definition 3.4 (Definition 5.1 of [14]).
Fix a field $\mathbb{k}$. For an oriented knot $K$ in $\mathbb{R}^{3}$, its augmentation variety over $\mathbb{k}$ is a subset of $(\mathbb{k}^{*})^{3}$ defined by
[…]
From Theorem 3.2, we obtain a refinement of [15, Theorem 1.1].
Corollary 3.5.
There exist $a\in\{\pm 1\}$ and integers $b,c$ and $d$ for which we have an injective map
$V_{\mathbb{k}}(K_{1})\hookrightarrow V_{\mathbb{k}}(K_{0})\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
@@
Using $\Psi_{\varphi}$ of Theorem 3.2, it follows that for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{k}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{k}$ such that
From the definition of $V_{\mathbb{k}}(K_{i})$ for $i\in\{0,1\}$, we obtain a map from $V_{\mathbb{k}}(K_{1})$ to $V_{\mathbb{k}}(K_{0})$ as in the assertion. ∎
The next proposition is used to prove Theorem 3.14 and Theorem 3.17 by taking $\mathbb{k}=\mathbb{Q}$. Consider a projection $\pi_{y,Z}\colon(\mathbb{k}^{*})^{3}\to(\mathbb{k}^{*})^{2}\colon(x,y,Z)\mapsto(y,Z)$.
@@
Suppose that $K_{1}$ is the unknot. Then,
$\{(y,Z)\in(\mathbb{k}^{*})^{2}\mid y\neq Z^{n}\text{ for all }n\in\mathbb{Z}\}\subset\pi_{y,Z}(V_{\mathbb{k}}(K_{0})).$
@@
If $K_{1}$ is the unknot, from the trivial braid on a single strand, we can compute as in [14, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by $U-\lambda-\mu+\lambda\mu$. Let us take $a\in\{\pm 1\}$ and $b,c,d\in\mathbb{Z}$ to be the integers in Corollary 3.5. For any $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, we take an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{k}$ determined by
@@
Note that $y\cdot Z^{-d}\neq 1$ and $y\cdot Z^{-d}\neq Z^{a}$ from the condition on $y$ and $Z$. By the map of Corollary 3.5, $(\varepsilon_{1}(\lambda),(y\cdot Z^{-d})^{a},Z)\in V_{\mathbb{k}}(K_{1})$ is sent to an element of $V_{\mathbb{k}}(K_{0})$ whose $(y,Z)$-component is $((y\cdot Z^{-d})\cdot Z^{d},Z)=(y,Z)$. ∎